OpenStudy (stylesprettygirl):

please algebra proof 8th grade? What is the correct reason for statement 3? Prove: (9m) ∙ 4 + m = 37m Statements Reasons 1. (9m) ∙ 4 + m 1. Given 2. 4 ∙ (9m) + m 2. Commutative Property 3. (4 ∙ 9) ∙ m + m 3. 4. 36m + m 4. Commutative Property 5. 37m 5. Addition A. Associative Property B. Multiplication C. Distributive Property D. Addition

OpenStudy (cmorris2004):

Hmm whats your guess...

OpenStudy (stylesprettygirl):

a

OpenStudy (anonymous):

Yes its a

OpenStudy (cmorris2004):

Well you are absolutely correct.. I think

OpenStudy (cmorris2004):

Hey this was my questin

OpenStudy (stylesprettygirl):

one more please i just need you guys to check

OpenStudy (cmorris2004):

-_-

OpenStudy (cmorris2004):

go on

OpenStudy (stylesprettygirl):

What is the correct reason for statement 5? Prove: k + 4 + 6(1 + k) = 11k + 10 Statements Reasons 1. k + 4 + 6(1 + k) 1. Given 2. k + 4 + (6+ 6k) 2. Distributive Property 3. k + (4 + 6) + 6k 3. Associative Property 4. k + 10 + 6k 4. Addition 5. k + 6k + 10 5. 6. 7k + 10 6. Addition A. Multiplication B. Commutative Property C. Associative Property D. Addition c i think?

OpenStudy (cmorris2004):

No. It is Multiplication or either Addition for 5 I think.

OpenStudy (cmorris2004):

brb

OpenStudy (stylesprettygirl):

i think it might be b

OpenStudy (cmorris2004):

OK i'm back

OpenStudy (cmorris2004):

Well go with your first thought it just my be right

OpenStudy (anonymous):

Oki here I am!!! and the CMorris dude is wrong. It's communitive property. Communitive Property: A+B=C B+A=C

OpenStudy (cmorris2004):

and where is your proof